OpenStudy (anonymous):

What mass of NaOH is in 240 mL of a 3.1 M NaOH solution?

OpenStudy (anonymous):

Molarity= Moles/ Volume in litres. 3.1=moles/0.24 moles=3.1*0.24=0.744. Weight=0.744*40=29.76